If a bag contains 6 red balls and 8 blue balls, and you pull one ball from the bag, which is blue, what is the probability that the next ball you pull will be red?


1
Expert's answer
2021-11-08T16:18:48-0500

Total number of balls is "=6+8=14balls"


Probability to pull one blue ball "=\frac {8}{14}
=\frac {4}{7}"


Making the total number of balls reduce from "14" to "13"

Hence

Probability that next ball is red "=\frac{6}{13}"